OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 3) and (5, –2)? begin fraction five over four negative begin fraction five over four begin fraction four over five negative begin fraction four over five

pooja195 (pooja195):

Use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{-2-3}{5-1}=slope\] Find the slope

OpenStudy (anonymous):

thx I get it now

pooja195 (pooja195):

yw ^_^